A simple pendulum has a period 2 second and is operated in a lift on earth 

what happen to its period 

i.) When the lift is ascending at uniform velocity

ii.) When the lift is ascending at an acceleration a m/s2

iii.) When the lift cable snaps?


1
Expert's answer
2021-03-24T19:38:59-0400

The period of a simple pendulum is:


"T = 2\pi\sqrt{\dfrac{l}{g}}"

where "g" is the acceleration of the pendulum and "l" is its length.

i. When lift is ascending at uniform velocity neither length nor acceleration changes, thus, the period remains the same.

ii. When the lift is ascending at an acceleration a m/s2 the resultant acceleration becomes a+g, and the period decreases.

iii. When the lift cable snaps, the acceleration becomes 0 and the pendulum does not oscillate.